OpenStudy (anonymous):

The floor of a room is 6m30cm long and 5m85cm wide. It is to be praded by square marble stone tiles. Find the largest size of each tile and hence find the least number of tiles required.

OpenStudy (anonymous):

let x be the side of the tile and y be the number of the tiles the area of the room = 6.3 * 5.85 = 36.855 the area of the tile = x^2 36.855 = x^2 y \[y = \frac{36.855}{x^2}\] draw this graph and see where y is smallest.